OpenStudy (yellowlegoguy99):

Please help with 8th grade math. Will give medals and fan.

OpenStudy (yellowlegoguy99):

Divide (56 x 10^7) / (7 x 10^21). Express the answer in scientific notation. 8 x 10^-14 8 x 10^-3 8 x 10^3 8 x 10^28

OpenStudy (yellowlegoguy99):

@Jack1

OpenStudy (yellowlegoguy99):

I went with A

OpenStudy (jack1):

\[\huge \frac {56 \times 10^7}{ 7 \times 10^{21}}\] \[\huge \frac {56 \times 10^{\color{red}{0}}}{ 7 \times 10^{(21 \color{red}{- 7})}}\] \[\huge \frac {56}{ 7 \times 10^{14}}\] \[\huge \frac {7 \times 8}{ 7 \times 10^{14}}\] \[\huge \frac { 8}{ 1 \times 10^{14}}\]

OpenStudy (jack1):

so yep, u were right, A

OpenStudy (yellowlegoguy99):

An easier way to do it would be, 56 / 7 = 8. 7 - 21 = -14. 8 x 10^-14

OpenStudy (jack1):

yep, wasn't sure if u wanted the steps though or just the answer dude